ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

15. Numerical Analysis and Scientific Computing

1991 MS Classification: 65, 68 Martínez, Raquel, Universidad Politécnica de Madrid, Spain Dupée, Brian, University of Bath, United Kingdom Tapia, Santiago, Universidad Politécnica de Madrid, Spain Composite Methods for Interpolating Functions The combination of the power of symbolic systems and the reliability of numerical methods has significant uses in the solution of mathematical problems. This presentation will describe a prototype package which uses algorithms created within the symbolic algebra system Axiom, numerical routines from the NAG Libraries together with the easy-to-use facilities of modern graphical interfaces. The application considers the implementation of different techniques, both symbolic and numeric, for the analysis and calculation of interpolating functions for curve fitting with the intention of providing scientists, engineers and their students with a graphical tool for the understanding of relationships between interpolation methods. This project was supported by Acciones Integradas Programme (HB1996-